Report on the Sabbatical Leave for the Full Academic Year 2022-2023

Doron ZEILBERGER, Board of Governors Professor, Department of Mathematics, Rutgers University

During the academic year 2022-2023 I was fortunate to have a sabbatical leave (80% salary). I used it to further my research activities, that cosiderably beneifited without the distaction of teaching. However I continued to regularly meet with my five PhD students via Zoom, and they made great progress. I also continued to organize my famous Zoom Experimental Mathematics Seminar.

Below is a list of the papers that were produced:

• Lots and Lots of Perrin-Type Primality Tests and Their Pseudo-Primes By Robert Dogherty-Bliss and Doron Zeilberger

• A Guide to the Risk-Averse Gambler and Resolving the St. Petersburg Paradox Once and For All By Lucy Martinez and Doron Zeilberger

• Experimenting with Discrete Dynamical System By George Spahn and Doron Zeilberger

• Using Symbolic Computation to Explore Generalized Dyck Paths and Their Areas By AJ Bu and Doron Zeiberger

For more information please consult my homepage

https://sites.math.rutgers.edu/~zeilberg/

Doron Zeilberger, Department of Mathematics, Rutgers University (New Brunswick), Hill Center-Busch Campus, 110 Frelinghuysen Rd., Piscataway, NJ 08854-8019, USA.

Email: zeilberg@math.rutgers.edu .

Sept. 27, 2023